# OmniCT: Towards a Unified Slice-Volume LVLM for Comprehensive CT Analysis

### v1

@@ 4 Experiments
[…]
4.1 Main Experiments
Slice-driven Understanding. We systematically evaluate OmniCT on four mainstream VQA benchmarks. As shown in the Table 4.1, medical LVLMs (e.g., HuatuoGPT-V-Qwen2.5, MedGemma) demonstrate relatively strong medical semantic understanding in certain tasks but remain limited in overall performance, often encountering bottlenecks on complex tasks. In contrast, general LVLMs achieve competitive or even superior results on some benchmarks, reflecting their strengths in language reasoning, but lack adaptation to CT images. For comparison, RadFM, although capable of handling both slice and volume inputs, achieves the weakest performance across all slice benchmarks, with an average score of only 32.12, failing to meet the demands of fine-grained CT tasks. Under the same evaluation protocol, OmniCT consistently surpasses existing models at both 3B and 7B scales: the 7B version achieves an average score of 81.45, exceeding the second-best model Lingshu by more than +11.01. These results demonstrate the robustness and comprehensiveness of OmniCT on slice-driven tasks.
@@ 4.1 Main Experiments
Volume-driven Understanding. As shown in Table 4.2, we further assess OmniCT on M3D, CT-RATE, and 3D-RAD to evaluate its volumetric perception capability for CT volume. Results show that existing volume-driven medical LVLMs (e.g., M3D-LaMed-7B/4B, CT-CHAT) achieve strong performance on specific subtasks—for example, CT-CHAT reaches 86.46 on CT-RATE multi-choice—but their overall averages remain below 36, highlighting limitations in coverage and stability. General LVLMs also exhibit strong cross-modal generalization in certain volume tasks, with GPT-5 leading multiple subtasks on 3D-RAD; however, their performance is highly inconsistent and lacks domain adaptation to CT volume. By contrast, OmniCT achieves clear advantages at both 3B and 7B scales: the 3B version reaches 87.38 on CT-RATE multi-choice with an average of 63.48, while the 7B version achieves 85.69 on the LTD task of 3D-RAD, pushing its overall average to 66.15—significantly outperforming all compared models. In addition, considering the significant importance of CT report generation, we performed 18-class abnormality label prediction for the report generation task on CT-RATE using fine-tuned RadBERT (Yan et al., 2022). The results show that OmniCT outperforms most volume-driven CT models and previous unified models (see Table 16), and performs similarly to models specifically designed for CT volume report generation (Hamamci et al., 2024b; Di Piazza et al., 2025). This validates the superiority of OmniCT in 3D spatial modeling and cross-task consistency. Across both slice-driven and volume-driven benchmarks, OmniCT demonstrates stable and comprehensive superiority at different scales, highlighting its holistic perception of spatial–semantic features in CT volume understanding tasks.
@@ 4.3 In-Depth Study
(ii) 2D Encoders vs. 3D Encoders. Given the inherent differences in design objectives and input modes, directly applying 3D encoders to 2D inputs often requires artificial adaptations such as depth replication, which compromises the fairness of comparison. Therefore, we conduct evaluations in native 3D settings. As shown in Figure 4 (b), even though M3D-CLIP (Bai et al., 2024) is pretrained with contrastive learning on the M3D dataset, it does not exhibit a clear advantage over 2D encoders such as DINOv3 (245 tokens) (Siméoni et al., 2025) and SigLIP (405 tokens) (Zhai et al., 2023), despite using the largest number of visual tokens (512). These results indicate that, at this stage, 2D encoders not only provide a more natural compatibility with both 2D and 3D inputs but also demonstrate stronger generalization across tasks, organs, and modalities. To assess the generality of this finding, we additionally evaluate several recent native 3D encoders (Wan et al., 2025; Wang et al., 2023) under the same protocol; the results are provided in Table 4. It is worth noting that we are not claiming that 2D features can fully represent 3D volumes. Instead, we offer a more measured assessment: at the current stage, through structured reorganization and volume-level embedding, more generalizable 2D encoders can robustly carry 3D spatial information. This design does not collapse dimensionality; rather, it retains spatial structures and relationships that remain interpretable from a 3D perspective on top of a 2D semantic backbone.
@@ 5 Related Work
Slice-driven Medical LVLMs. With the success of Large Language Models in general language (Liu et al., 2024; Yang et al., 2025) and vision (Qiu et al., 2020; Bai et al., 2025a), early explorations focused on adapting general LVLM paradigms to the medical domain, such as LLaVA-Med (Li et al., 2023) and Med-Flamingo (Moor et al., 2023), which leveraged medical image–text pairs and instruction data to enable initial medical capabilities. Subsequently, a series of more general-purpose medical LVLMs emerged, including RadFM (Wu et al., 2025), BiomedGPT (Luo et al., 2023), HuatuoGPT-Vision (Chen et al., 2024a), and Lingshu (Xu et al., 2025). These models advanced the field through large-scale data curation (Bansal et al., 2024), reasoning-enhanced training strategies (Pan et al., 2025; Xu et al., 2025), multi-task generalization (Jiang et al., 2024), and domain-specific knowledge integration (Sellergren et al., 2025). Recently, models such as CXR-LLaVA (Lee et al., 2025) and EyecareGPT (Li et al., 2025a) have demonstrated stronger adaptability and diagnostic value in modality-specific and specialty-oriented tasks (Xie et al., 2025; Hao et al., 2025). Nevertheless, despite substantial progress in data scale, architectural design, and task diversity, slice-driven medical LVLMs remain constrained by their reliance on planar inputs, limiting their ability to capture the spatial consistency and cross-slice dependencies essential for CT understanding.
Volume-driven Medical LVLMs. To overcome the limitations of 2D modeling, research has increasingly turned to 3D volumetric imaging, employing dedicated datasets, 3D encoders, and cross-modal alignment modules to strengthen spatial modeling in clinical tasks (Wu et al., 2025). M3D-LaMed (Bai et al., 2024) established a comprehensive evaluation system across multiple volumetric medical tasks, while CT-CHAT (Hamamci et al., 2024c) introduced paired chest CT data and an architecture tailored for fine-grained analysis and dialog-based interaction. At the methodological level, Med-2E3 (Shi et al., 2024) combined 2D and 3D encoders and enhanced reasoning consistency through dynamic cross-slice scoring, whereas Med3DInsight (Chen et al., 2024c) aligned a 3D encoder with a 2D LVLM, achieving strong performance in both segmentation and classification. Nevertheless, the lack of a unified clinical evaluation framework and efficient slice–volume collaboration mechanisms continues to limit adaptability and scalability.
@@
[…]
@@ 6 Conclusion
Ethics statement
This work adheres to the ICLR Code of Ethics. No human subjects or animal experimentation were involved in this study. All medical imaging datasets used in our experiments, including SLAKE, VQA-RAD, OmniMedVQA, RadFig-VQA, M3D, CT-RATE, and 3D-RAD, are publicly available under relevant research licenses and comply with usage guidelines, ensuring no violation of privacy or ethical standards. No personally identifiable information (PII) was included, and all CT data were pre-processed (e.g., windowing, resampling, and anonymization) before use. We took special care to avoid any misuse of data and to ensure that all evaluations were conducted fairly and transparently. Our contributions focus solely on methodological and benchmarking innovations, without clinical or diagnostic decision-making implications.
Reproducibility Statement
We have made every effort to ensure the reproducibility of our results. All codes and datasets will be released in an open repository with detailed documentation. The experimental setup, including model architectures, training hyperparameters, optimization strategies, is described in full detail in the main text and appendix. In addition, we rely on multiple widely used public datasets (e.g., SLAKE, VQA-RAD, M3D, CT-RATE) to facilitate verification and cross-comparison. Our proposed Data Orchestration Engine ensures consistent data preprocessing and evaluation, further improving reproducibility across different models. We believe these measures will enable the community to replicate our work, benchmark future models fairly, and extend the development of medical LVLMs in CT understanding.
@@ Appendix C Notation Table
[…]
@@
[…]
@@
[…]
@@ Appendix D Extended Related Work
In recent times, the release of multiple Multi-modal Large Language Models (MLLMs) has driven innovations in vision-language fusion, long temporal sequence processing, and scenario adaptability (Hong et al., 2025; Team et al., 2025), laying a solid foundation for cross-domain applications. Prominent foundation models such as Qwen2.5-VL (Bai et al., 2025b), GPT-4o (Achiam et al., 2023), Claude 3.5 (AI, 2024), InternVL3 (Zhu et al., 2025), and the latest GPT-5 (Wang et al., 2025) have continuously advanced in multi-modal understanding, long-sequence processing, multi-task learning, and vertical domains like healthcare (Arora et al., 2025), demonstrating exceptional potential. These advancements are primarily driven by high-quality data curation and iterative algorithmic optimization. However, as foundation models, maintaining a balance between general-purpose capabilities and domain-specific expertise remains a significant challenge.
@@ Appendix E Implementation Details
Data Details. For 2D slice evaluation, we construct test sets based on SLAKE (Liu et al., 2021), VQA-RAD (Lau et al., 2018), OmniMedVQA (Hu et al., 2024), and RadFig (Yamagishi et al., 2025), where all samples are systematically filtered by the data engine (Section 3) to retain only high-quality CT VQA data. For 3D volume evaluation, we adopt existing benchmarks including M3D (Bai et al., 2024), CT-RATE (Hamamci et al., 2024a), and 3D-RAD (Gai et al., 2025) to cover the full spectrum of CT volumetric scenarios. Regarding evaluation metrics, Accuracy is used for closed-end and multiple-choice tasks, while open-ended QA is assessed by a weighted combination of BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), Token-F1 (Saab et al., 2024), and BERTScore (Zhang et al., 2019), balancing lexical matching with semantic alignment to achieve multi-level quality measurement. For data pre-processing, all CT volumes with preserved metadata are windowed to [1000,1000][-1000,1000] and resampled to $32\times 384\times 384$.
Model Details and Implementation Details. We use siglip-so400m-patch14-384 (Zhai et al., 2023) as the vision encoder and Qwen2.5 (Team, 2024) as the backbone LLM, with AdamW (Loshchilov and Hutter, 2017) as the optimizer. During pretraining, only the MoE Hybrid Projection is updated to perform cross-modal alignment, with a learning rate of $2\times 10^{-4}$. In the vision instruction tuning stage, both the projection layer and LLM parameters are optimized, with the learning rate reduced to $5\times 10^{-5}$ to ensure stable convergence. All experiments are trained with a global batch size of $256$ and a warmup–cosine learning rate scheduler. Unless otherwise specified, experiments are conducted under the 7B parameter scale. The specific hyperparameter settings can be found in Table 6.
@@ Appendix F Data Orchestration Engine
The construction of MedEval-CT is powered by the proposed Data Orchestration Engine, which integrates four complementary modules. Specifically, the Corpus Selector and Integrity Verifier are implemented with Qwen2.5-VL-72B (Bai et al., 2025b), while the Task Mapper and Semantic Refiner leverage Qwen3-237B-A3B (Yang et al., 2025), thereby exploiting the complementary strengths of different models in large-scale data filtering and semantic refinement. The specific prompt designs for each module can be found in Fig. 5.
To ensure the reliability and independence of MedEval-CT, we proactively conduct a systematic audit of potential data overlap during its construction. For the slice datasets, we note that ROCOv2 (Ronan L.M., 2024), PubMedVision (Chen et al., 2024b), LLaVA-Med (Li et al., 2023), and RadFig-VQA (Yamagishi et al., 2025) originate from PMC-OA (Lin et al., 2023b). Although these four datasets follow their own automated curation pipelines, we further apply a two-stage deduplication strategy within them: perceptual image hashing (pHash) (Monga and Evans, 2006) is used to cluster visually similar images, followed by BiomedCLIP (Zhang et al., 2023) feature matching to remove image–text pairs with high semantic similarity. For the volume datasets, we strictly adhere to the official splits of M3D (Bai et al., 2024), CT-RATE (Hamamci et al., 2024a), and 3D-RAD (Gai et al., 2025); even in cases where datasets share underlying CT volumes, we avoid any cross-dataset training or evaluation. Throughout the pipeline, we retain only modality-consistent and high-quality CT scans, filtering out blurry, artifact-heavy, or low-resolution samples. These measures allow MedEval-CT to maintain strict separation in data sourcing, partitioning, and deduplication, effectively minimizing the risks of training–testing contamination and data leakage.
To further address the concern that using Qwen-family models in both the data pipeline and the LLM base model might introduce family-specific bias or circularity, we additionally instantiate OmniCT with a different LLM backbone, Phi-4-mini (Abouelenin et al., 2025), while keeping the training data and optimization protocol unchanged. As shown in Table 7, OmniCT with Phi-4-mini achieves performance that is highly comparable to, and on several benchmarks slightly better than, the Qwen2.5-3B variant across both slice-driven (SLAKE, VQA-RAD, OmniMedVQA, RadFig-VQA) and volume-driven (M3D, CT-RATE, 3D-RAD, MedEval-CT-Bench) benchmarks. This consistency indicates that the observed gains mainly stem from the proposed unified framework itself rather than from any base model-specific preference or bias induced by the models used in the data construction pipeline.
@@ Appendix G Mechanism of MedEval-CT-Factory
Multi-dimensional Evaluation Protocols. At the output level, MedEval-CT-Factory integrates multiple evaluation strategies to accommodate the diversity of outputs produced by medical LVLMs. Rather than imposing a rigid scoring pipeline, it provides a composable and extensible evaluation space: (i) statistical metrics (BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005)) for measuring surface-level textual alignment; (ii) semantic metrics (BERTScore (Zhang et al., 2019), embedding similarity (Zhang et al., 2025b)) for assessing semantic correspondence; and (iii) LLM-based evaluation for simulating clinical reasoning, offering more qualitative judgments aligned with medical scenarios. Users may flexibly select appropriate evaluation layers according to task requirements without being restricted to a single metric.

### v2

@@ 4.1 Main Experiments
81.45 66.15 78.24 80.27 79.62
Slice-driven Understanding. We systematically evaluate OmniCT on four mainstream VQA benchmarks. As shown in the Table 2, medical LVLMs (e.g., HuatuoGPT-V-Qwen2.5, MedGemma) demonstrate relatively strong medical semantic understanding in certain tasks but remain limited in overall performance, often encountering bottlenecks on complex tasks. In contrast, general LVLMs achieve competitive or even superior results on some benchmarks, reflecting their strengths in language reasoning, but lack adaptation to CT images. For comparison, RadFM, although capable of handling both slice and volume inputs, achieves the weakest performance across all slice benchmarks, with an average score of only 32.12, failing to meet the demands of fine-grained CT tasks. Under the same evaluation protocol, OmniCT consistently surpasses existing models at both 3B and 7B scales: the 7B version achieves an average score of 81.45, exceeding the second-best model Lingshu by more than +11.01. These results demonstrate the robustness and comprehensiveness of OmniCT on slice-driven tasks.
@@
Volume-driven Understanding. As shown in Table 3, we further assess OmniCT on M3D, CT-RATE, and 3D-RAD to evaluate its volumetric perception capability for CT volume. Results show that existing volume-driven medical LVLMs (e.g., M3D-LaMed-7B/4B, CT-CHAT) achieve strong performance on specific subtasks—for example, CT-CHAT reaches 86.46 on CT-RATE multi-choice—but their overall averages remain below 36, highlighting limitations in coverage and stability. General LVLMs also exhibit strong cross-modal generalization in certain volume tasks, with GPT-5 leading multiple subtasks on 3D-RAD; however, their performance is highly inconsistent and lacks domain adaptation to CT volume. By contrast, OmniCT achieves clear advantages at both 3B and 7B scales: the 3B version reaches 87.38 on CT-RATE multi-choice with an average of 63.48, while the 7B version achieves 85.69 on the LTD task of 3D-RAD, pushing its overall average to 66.15—significantly outperforming all compared models. In addition, considering the significant importance of CT report generation, we performed 18-class abnormality label prediction for the report generation task on CT-RATE using fine-tuned RadBERT (yan2022radbert). The results show that OmniCT outperforms most volume-driven CT models and previous unified models (see Table 16), and performs similarly to models specifically designed for CT volume report generation (hamamci2024ct2rep; di2025ct). This validates the superiority of OmniCT in 3D spatial modeling and cross-task consistency. Across both slice-driven and volume-driven benchmarks, OmniCT demonstrates stable and comprehensive superiority at different scales, highlighting its holistic perception of spatial–semantic features in CT volume understanding tasks.
@@ 4.3 In-Depth Study
(ii) 2D Encoders vs. 3D Encoders. Given the inherent differences in design objectives and input modes, directly applying 3D encoders to 2D inputs often requires artificial adaptations such as depth replication, which compromises the fairness of comparison. Therefore, we conduct evaluations in native 3D settings. As shown in Figure 4 (b), even though M3D-CLIP (bai2024m3d) is pretrained with contrastive learning on the M3D dataset, it does not exhibit a clear advantage over 2D encoders such as DINOv3 (245 tokens) (simeoni2025dinov3) and SigLIP (405 tokens) (zhai2023sigmoid), despite using the largest number of visual tokens (512). These results indicate that, at this stage, 2D encoders not only provide a more natural compatibility with both 2D and 3D inputs but also demonstrate stronger generalization across tasks, organs, and modalities. To assess the generality of this finding, we additionally evaluate several recent native 3D encoders (wan2025wan; wang2023videomaev2) under the same protocol; the results are provided in Table 4. It is worth noting that we are not claiming that 2D features can fully represent 3D volumes. Instead, we offer a more measured assessment: at the current stage, through structured reorganization and volume-level embedding, more generalizable 2D encoders can robustly carry 3D spatial information. This design does not collapse dimensionality; rather, it retains spatial structures and relationships that remain interpretable from a 3D perspective on top of a 2D semantic backbone.
@@ 5 Related Work
Encoder Token Budget Ratio Plane Phase Organ Abnormality Location Avg.
@@
Slice-driven Medical LVLMs. With the success of Large Language Models in general language (liu2024deepseek; yang2025qwen3) and vision (qiu2020dgcn; Qwen3-VL), early explorations focused on adapting general LVLM paradigms to the medical domain, such as LLaVA-Med (li2023llava) and Med-Flamingo (moor2023med), which leveraged medical image–text pairs and instruction data to enable initial medical capabilities. Subsequently, a series of more general-purpose medical LVLMs emerged, including RadFM (wu2025towards), BiomedGPT (luo2023biomedgpt), HuatuoGPT-Vision (chen2024huatuogpt), and Lingshu (xu2025lingshu). These models advanced the field through large-scale data curation (bansal2024medmax), reasoning-enhanced training strategies (pan2025medvlm; xu2025lingshu), multi-task generalization (jiang2024med), and domain-specific knowledge integration (sellergren2025medgemma). Recently, models such as CXR-LLaVA (lee2025cxr) and EyecareGPT (li2025eyecaregpt) have demonstrated stronger adaptability and diagnostic value in modality-specific and specialty-oriented tasks (xie2025heartcare; hao2025towards). Nevertheless, despite substantial progress in data scale, architectural design, and task diversity, slice-driven medical LVLMs remain constrained by their reliance on planar inputs, limiting their ability to capture the spatial consistency and cross-slice dependencies essential for CT understanding.
Volume-driven Medical LVLMs. To overcome the limitations of 2D modeling, research has increasingly turned to 3D volumetric imaging, employing dedicated datasets, 3D encoders, and cross-modal alignment modules to strengthen spatial modeling in clinical tasks (wu2025towards). M3D-LaMed (bai2024m3d) established a comprehensive evaluation system across multiple volumetric medical tasks, while CT-CHAT (hamamci2024developing) introduced paired chest CT data and an architecture tailored for fine-grained analysis and dialog-based interaction. At the methodological level, Med-2E3 (shi2024med) combined 2D and 3D encoders and enhanced reasoning consistency through dynamic cross-slice scoring, whereas Med3DInsight (chen2024med3dinsight) aligned a 3D encoder with a 2D LVLM, achieving strong performance in both segmentation and classification. Nevertheless, the lack of a unified clinical evaluation framework and efficient slice–volume collaboration mechanisms continues to limit adaptability and scalability.
@@ 6 Conclusion
Acknowledgement
This work has been supported in part by the NSFC (No. 62436007), the China Postdoctoral Science Foundation under Grant Number 2024M752794, the ZJNSF (No. LZ25F020004), the Key Research and Development Projects in Zhejiang Province (No. 2025C01128, 2025C01030, 2025C02156), Zhejiang University Education Foundation Qizhen Scholar Foundation.
@@ Appendix C Notation Table
OmniCT-3B OmniCT-7B
Hyperparameter Stage-1 Stage-2 Stage-1 Stage-2
@@
Base Model #Params SLAKE VQA-RAD OmniMed VQA RadFig -VQA M3D CT-RATE 3D-RAD MedEval-CT -Bench
@@
Dataset Task Type Train Size Test Size Source License
Volume-driven Datasets
@@ Appendix D Extended Related Work
In recent times, the release of multiple Multi-modal Large Language Models (MLLMs) has driven innovations in vision-language fusion, long temporal sequence processing, and scenario adaptability (hong2025glm; team2025gemma), laying a solid foundation for cross-domain applications. Prominent foundation models such as Qwen2.5-VL (bai2025qwen2), GPT-4o (achiam2023gpt), Claude 3.5 (TheC3), InternVL3 (zhu2025internvl3), and the latest GPT-5 (wang2025capabilities) have continuously advanced in multi-modal understanding, long-sequence processing, multi-task learning, and vertical domains like healthcare (arora2025healthbench), demonstrating exceptional potential. These advancements are primarily driven by high-quality data curation and iterative algorithmic optimization. However, as foundation models, maintaining a balance between general-purpose capabilities and domain-specific expertise remains a significant challenge.
@@ Appendix E Implementation Details
Data Details. For 2D slice evaluation, we construct test sets based on SLAKE (liu2021slake), VQA-RAD (lau2018dataset), OmniMedVQA (hu2024omnimedvqa), and RadFig (yamagishi2025radfig), where all samples are systematically filtered by the data engine (Section 3) to retain only high-quality CT VQA data. For 3D volume evaluation, we adopt existing benchmarks including M3D (bai2024m3d), CT-RATE (hamamci2024foundation), and 3D-RAD (gai20253d) to cover the full spectrum of CT volumetric scenarios. Regarding evaluation metrics, Accuracy is used for closed-end and multiple-choice tasks, while open-ended QA is assessed by a weighted combination of BLEU (papineni2002bleu), ROUGE (lin2004rouge), Token-F1 (saab2024capabilities), and BERTScore (zhang2019bertscore), balancing lexical matching with semantic alignment to achieve multi-level quality measurement. For data pre-processing, all CT volumes with preserved metadata are windowed to [1000,1000][-1000,1000] and resampled to $32\times 384\times 384$.
Model Details and Implementation Details. We use siglip-so400m-patch14-384 (zhai2023sigmoid) as the vision encoder and Qwen2.5 (team2024qwen2) as the backbone LLM, with AdamW (loshchilov2017decoupled) as the optimizer. During pretraining, only the MoE Hybrid Projection is updated to perform cross-modal alignment, with a learning rate of $2\times 10^{-4}$. In the vision instruction tuning stage, both the projection layer and LLM parameters are optimized, with the learning rate reduced to $5\times 10^{-5}$ to ensure stable convergence. All experiments are trained with a global batch size of $256$ and a warmup–cosine learning rate scheduler. Unless otherwise specified, experiments are conducted under the 7B parameter scale. The specific hyperparameter settings can be found in Table 6.
@@ Appendix F Data Orchestration Engine
The construction of MedEval-CT is powered by the proposed Data Orchestration Engine, which integrates four complementary modules. Specifically, the Corpus Selector and Integrity Verifier are implemented with Qwen2.5-VL-72B (bai2025qwen2), while the Task Mapper and Semantic Refiner leverage Qwen3-237B-A3B (yang2025qwen3), thereby exploiting the complementary strengths of different models in large-scale data filtering and semantic refinement. The specific prompt designs for each module can be found in Fig. 5.
To ensure the reliability and independence of MedEval-CT, we proactively conduct a systematic audit of potential data overlap during its construction. For the slice datasets, we note that ROCOv2 (ronan_l.m._2024), PubMedVision (chen2024huatuogptvisioninjectingmedicalvisual), LLaVA-Med (li2023llava), and RadFig-VQA (yamagishi2025radfig) originate from PMC-OA (lin2023pmc). Although these four datasets follow their own automated curation pipelines, we further apply a two-stage deduplication strategy within them: perceptual image hashing (pHash) (monga2006perceptual) is used to cluster visually similar images, followed by BiomedCLIP (zhang2023biomedclip) feature matching to remove image–text pairs with high semantic similarity. For the volume datasets, we strictly adhere to the official splits of M3D (bai2024m3d), CT-RATE (hamamci2024foundation), and 3D-RAD (gai20253d); even in cases where datasets share underlying CT volumes, we avoid any cross-dataset training or evaluation. Throughout the pipeline, we retain only modality-consistent and high-quality CT scans, filtering out blurry, artifact-heavy, or low-resolution samples. These measures allow MedEval-CT to maintain strict separation in data sourcing, partitioning, and deduplication, effectively minimizing the risks of training–testing contamination and data leakage.
To further address the concern that using Qwen-family models in both the data pipeline and the LLM base model might introduce family-specific bias or circularity, we additionally instantiate OmniCT with a different LLM backbone, Phi-4-mini (abouelenin2025phi), while keeping the training data and optimization protocol unchanged. As shown in Table 7, OmniCT with Phi-4-mini achieves performance that is highly comparable to, and on several benchmarks slightly better than, the Qwen2.5-3B variant across both slice-driven (SLAKE, VQA-RAD, OmniMedVQA, RadFig-VQA) and volume-driven (M3D, CT-RATE, 3D-RAD, MedEval-CT-Bench) benchmarks. This consistency indicates that the observed gains mainly stem from the proposed unified framework itself rather than from any base model-specific preference or bias induced by the models used in the data construction pipeline.
@@ Appendix G Mechanism of MedEval-CT-Factory
Multi-dimensional Evaluation Protocols. At the output level, MedEval-CT-Factory integrates multiple evaluation strategies to accommodate the diversity of outputs produced by medical LVLMs. Rather than imposing a rigid scoring pipeline, it provides a composable and extensible evaluation space: (i) statistical metrics (BLEU (papineni2002bleu), ROUGE (lin2004rouge), METEOR (banerjee2005meteor)) for measuring surface-level textual alignment; (ii) semantic metrics (BERTScore (zhang2019bertscore), embedding similarity (qwen3embedding)) for assessing semantic correspondence; and (iii) LLM-based evaluation for simulating clinical reasoning, offering more qualitative judgments aligned with medical scenarios. Users may flexibly select appropriate evaluation layers according to task requirements without being restricted to a single metric.